K-types of minimal representations (p-adic case)

Gordan Savin, Salt Lake City, USA

AMS Subject classification: 22E35, 22E50, 11F70

Abstract: Let F be a p-adic field. Let G be a split simple simply connected group over F of type D_{n}, (n ≥ 4), or E_{n}, (n = 6, 7, 8). Let K be a hyperspecial maximal compact subgroup of G. In this article we describe K-types of the minimal representation of G.